OpenStudy (anonymous):

Use implicit differentiation to find y' and evaluate y' at the indicated point. x^2y-3x^2-4=0; (2,4)

OpenStudy (anonymous):

x^2y' + y(2x) - 6x = 0 x^2y' = 6x - 2xy y' = (6x - 2xy) / x^2 = [6(2) - 2(2)(4)] / 2^2 = (12 - 16) / 4 = -1